# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ 4 Application in the sequential filter
We call the regularized Ensemble Kalman Filter (EnKF) the Locally Stationary Ensemble Filter (LSEF). Its forecast step coincides with that of the stochastic EnKF. The only difference between the LSEF and the EnKF analysis is how the proxy $\widehat{\bf B}$ to the true prior covariance matrix is defined. In this way, we can isolate the impact of our spatial-model-based regularization approach.
@@
In the LSEF analysis scheme, we postulate that the forecast error obeys the Locally Stationary Convolution Model. The model is estimated from the forecast ensemble (as detailed in section 3) and used in the analysis to compute the analysis gain matrix. Specifically, having the estimate $\widehat{u}(x,\rho(x,y))$ of the true spatial convolution kernel $u(x,\rho(x,y))$, we build the space-discrete version of Eq. (3):
@@
where $\boldsymbol{\xi}$ is the vector that comprises the random field $\xi(x)$ evaluated at the points of a spatial grid, $\boldsymbol{\alpha}\sim\mathcal{N}({\bf 0,I})$ and $\widehat{\bf W}$ is a matrix with the entries
$\widehat{w}_{ij}=\widehat{u}(x_{i},\rho(x_{i},y_{j}))\,\sqrt{\Delta y_{j}},$ (16)
where $\Delta y_{j}$ is the area of $j$th grid cell. Note that Eqs. (15) and (16) approximate the integral in Eq. (3) using the simplest rectangle rule, more accurate approximations will be considered elsewhere. Equation (15) implies that
@@
We note that the representation of the forecast-error covariance matrix $\widehat{\bf B}$ in the “square-root” form, Eq. (17), is common in data assimilation practice because it provides efficient preconditioning of the analysis equations, e.g., Asch et al. (2016). The need for preconditioning follows from Eq. (29) because the matrix ${\bf B}^{-1}+{\bf H}^{\textsf{T}}{\bf R}^{-1}{\bf H}$ to be inverted is normally ill-conditioned. Substituting Eq. (17) into Eq. (29) and rearranging the terms yields
@@
Now the matrix to be inverted, ${\bf I}+\widehat{\bf W}^{\textsf{T}}{\bf H}^{\textsf{T}}{\bf R}^{-1}{\bf H}\widehat{\bf W}$, is, clearly, well conditioned (as a sum of the identity matrix and a positive definite matrix). This is how the LSEF analysis is done, that is, without explicitly computing $\widehat{\bf B}$. The LSEF analysis assimilates all available observations at once, as this is done in variational data assimilation schemes (e.g., Asch et al., 2016).
@@ 5 Numerical experiments with static analyses
In this section, we experimentally examine the above LSEF analysis technique in the setting with known synthetic “truth” that obeys the Locally Stationary Convolution Model on the sphere. This setting implies that the data generation model coincides with the model we assume in the LSEF analysis, which allowed us to reveal the best possible performance our approach can provide.
We compared deterministic analyses produced by the following schemes: (i) the optimal (Kalman filter) analysis which had access to the true prior covariance matrix ${\bf B}$, (ii) the regularized-EnKF (referred to as LSEF) analysis, (iii) the traditional-EnKF (referred to as just EnKF in the sequel) analysis, (iv) the analysis that relies on a static covariance matrix, and (v) a hybrid of the two latter analyses. We will refer to the five analyses as True-B, LSEF-B, EnKF-B, Mean-B, and Hybrid-B, respectively. See section 4 for the description of the LSEF analysis and A.3 and A.4 for the description of the other (reference) analyses.
The analysis ensemble is not tested here, its performance will be indirectly assessed in sequential filtering experiments presented in section 6.
@@ 5.1 Experimental methodology
Pseudo-random realizations of the “true” local spectra $f_{\ell}(x)$ and the respective “true” background error field $\xi(x)$ were generated following the Parametric Locally Stationary Convolution Model of “truth” described in J. The true background-error vector $\boldsymbol{\xi}$ was computed by evaluating $\xi(x)$ on the analysis grid (defined below in section 5.2). We assumed (without loss of generality) that ${\bf x}^{\rm f}=0$ so that the simulated truth was ${\bf x}={\bf x}^{\rm f}-\boldsymbol{\xi}=-\boldsymbol{\xi}$ and the ensemble members were ${\bf x}^{{\rm fe}(\mu)}={\bf x}^{\rm f}-\boldsymbol{\xi}^{(\mu)}=-\boldsymbol{\xi}^{(\mu)}$ (see A for notation).
@@
Having computed a realization of $f_{\ell}(x)$ at all spatial grid points, we computed $\sigma_{\ell}(x)=\sqrt{f_{\ell}(x)}$ and made use of Eq. (4) to compute the true kernel $u(x,\rho)$. After that, we built the true ${\bf W}$ matrix following Eq. (16) and applied it to realizations of the white noise to generate the background-error vector $\boldsymbol{\xi}$ and the ensemble perturbations $\boldsymbol{\xi}^{(\mu)}$, see Eq. (15). The true covariance matrix (used by the optimal “True-B” analysis) was ${\bf B}={\bf W}{\bf W}^{\textsf{T}}$.
@@
With the truth ${\bf x}$ in hand, we generated observations ${\bf x}^{\rm obs}$ following Eq. (25), assimilated them using the five analysis techniques listed above, and compared the resulting deterministic analyses with the truth.
@@
The bandpass filters’ transfer functions were specified as follows:
$H_{j}(\ell)=\exp\left(-\left|\frac{\ell-\ell^{\rm c}_{j}}{\Delta_{H_{j}}}\right|^{q_{\rm shape}}\right),$ (19)
@@
where $\ell^{\rm c}_{j}$ is the central wavenumber of the $j$th filter, $\Delta_{H_{j}}$ is the the half-width parameter, and $q_{\rm shape}$ is the shape parameter. In the experiments described below, we took ${q_{\rm shape}}$ equal to 3 (${q_{\rm shape}}=2$ also worked well). Both $\ell^{\rm c}_{j}$ and ${\Delta_{H_{j}}}$ were taken exponentially growing with the band index $j$. Figure 1 shows an example of the spectral transfer functions and the respective impulse response functions for six bandpass filters used in the experiments presented in section 5. The spatial filtering was performed in spectral space.
@@
The four sub-optimal analysis schemes differed only in the way the prior covariance matrix was specified, see A.3, A.4, and section 4. In the EnKF analysis, the localization matrix was specified using the popular Gaspari-Cohn correlation function (Gaspari and Cohn, 1999), whose length scale was tuned to obtain the best performance for each configuration of the model of truth and for each ensemble size separately.
In the “Mean-B” analysis, the static prior covariance matrix ${\bf B}_{\rm static}={\bf B}_{\rm mean}$ was computed for each configuration of the model of truth as follows. We generated $n_{\rm repl}=33$ replicates of the true local spectra $f_{\ell}(x)$. For each replicate, given $f_{\ell}(x)$, we generated a sample of $M=10$ realizations of random fields $\xi^{(\mu)}(x)$ (see above in this subsection). Then, we expanded each $\xi^{(\mu)}(x)$ in spherical harmonics, getting the spectral coefficients $\widetilde{\xi}^{(\mu)}_{\ell m}$. For each $\ell$, we summed $|\widetilde{\xi}^{(\mu)}_{\ell m}|^{2}$ over $m$ in the range from $-\ell$ to $\ell$, divided by $4\pi$, and averaged over $\mu=1,2,\dots,M$ and $n_{\rm repl}$ replicates of $f_{\ell}(x)$, obtaining a sample estimate of the mean variance spectrum of $\xi$. From the mean variance spectrum, we computed the mean (stationary in space) covariance function used to build ${\bf B}_{\rm mean}$.
In the “Hybrid-B” analysis, a convex combination of the static and localized sample covariance matrices was used: $\widehat{\bf B}_{\rm hybr}=(1-w_{\rm e}){\bf B}_{\rm mean}+w_{\rm e}\widehat{\bf B}_{\rm EnKF}$, where $10<w_{\rm e}<$ 1 is the weight to be tuned. $w_{\rm e}=0.5$ was found to be nearly optimal and used in all experiments.
@@ 5.2 Experimental setup
The default setup of the parametric model of truth is specified in J.4.
@@
Observations were located at randomly and independently sampled grid points selected with probabilities proportional to the spherical surfaces of the respective grid cells. The number of observations was $n_{\rm obs}=\frac{1}{2}n_{\rm x}$ and the observation-error variance was equal to the median background-error variance (we tried other choices and obtained qualitatively very similar results, not shown). Recall that $n_{\rm x}$ stands for the total number of grid points (the length of the state vector ${\bf x}$, see A).
The ensemble size was between 5 and 80 with the default 20. The number of bandpass filters was 6.
@@ 5.3 Training sample
Training the neural network involved in the estimation of the Locally Stationary Convolution Model was done as follows. The Parametric Locally Stationary Convolution Model, see J, was used to generate $n_{\rm repl}=33$ replicates the true local spectra $f_{\ell}(x)$. For each replicate of $\sigma_{\ell}(x)=\sqrt{f_{\ell}(x)}$, we generated the ensemble $\{\xi^{(\mu)}(x)\}_{\mu=1}^{M}$ and computed the set of sample band variances $d_{j}(x)$ at each analysis grid point $x$ (sections 3.2 and 5.1). As a result, with the spatial analysis grid with $n_{\rm x}$ points, we had the training sample of $n_{\rm train}=n_{\rm repl}\cdot n_{\rm x}$ pairs of vectors. The first vector in a pair (input of the neural network) consisted of $J$ sample band variances at the grid point $x$. The second vector (of length $\ell_{\rm max}+1$) in the pair (target or “truth”) was the respective “true” spectral function $\sigma_{\ell}(x)$.
@@ 5.4 Results
We compared the five deterministic analyses in the default experimental setup and examined effects of varying the ensemble size $M$ and the hyperparameters $\varkappa$ (controls the strength of the non-stationarity, J.2) and $\mu_{\rm NSL}$ (controls the length scale of the non-stationarity, J.3).
Figure 2 demonstrates the role of the ensemble size. It is seen that all ensemble analyses improved with the increasing $M$, as it should be. The advantage of the LSEF-B analysis over the competitors was very significant. Compared with the EnKF-B, the LSEF-B analysis performed especially well for small $M$. This is because using the Locally Stationary Convolution Model to compute the prior covariance matrix in LSEF-B effectively reduces sampling noise (from which EnKF suffers) but at the expense of introducing a systematic distortion in the prior covariances. In stronger sampling noise (i.e., with a smaller $M$), the systematic error is relatively less important, therefore, the advantage of LSEF-B is more pronounced with small ensembles.
@@
Figure 3 shows how the five analyses performed with the varying non-stationarity strength parameter $\varkappa$. The LSEF-B analysis’ advantage over EnKF-B was big and nearly uniform for all $\varkappa>1$ on the plot. It was Mean-B that suffered/benefited most from strong/weak non-stationarity, as expected. In the stationary regime (i.e., with $\varkappa=1$), Mean-B was almost as accurate as True-B because ${\bf B}_{\rm mean}$ was estimated by averaging over 33 replicates of the truth (in which the true local spectra were the same, see J.5) and also by spatial averaging (section 5.1). For this reason, Mean-B could not be rivaled by any ensemble analysis in the stationary regime.
@@
Figure 4 shows how the analyses performed with the varying non-stationarity length parameter $\mu_{\rm NSL}$. Again, LSEF-B was significantly better than the other sub-optimal analyses and that was the case even beyond the local stationarity assumption (section 2.6), which requires that the non-stationarity length scale is much larger than the typical length scale of the field itself. Indeed, LSEF-B was substantially better than the competing techniques even when $\mu_{\rm NSL}=1$ (see Fig. 4), that is, when the non-stationarity length scale was equal to the median local length scale (see Eqs. (85) and (82)).
@@
It is important to notice that in LSEF-B, the neural network was trained only once, for the default setup, whereas we had to tune the localization length scale in EnKF-B for each triple $M,\varkappa,\mu_{\rm NSL}$. Similarly, ${\bf B}_{\rm mean}$ was computed for use in Mean-B and Hybrid-B for every pair $\varkappa,\mu_{\rm NSL}$. This suggests that the neural net-based LSEF-B was more robust than the competing analyses.
The neural network was trained on the sample of the same size as we used to calculate ${\bf B}_{\rm mean}$, yet it allowed LSEF-B to exhibit much better results than Hybrid-B. This suggests that the neural network is capable of selective blending of sample covariances with time-mean covariances. This is in contrast to Hybrid-B, which just mixes the two kinds of covariances with, effectively, equal weights for all spatial scales.
5.5 Effects of covariance regularization
We reproduced the above experimental methodology (the model of “truth” and the five analyses) on the circle and obtained very similar results (not shown). The code was written by the other author in a different programming language (R vs. Python).
Here we take advantage of the simpler circular domain to demonstrate how the regularization procedure proposed in this paper works. The spatial grid had 120 points and the ensemble size was 10.
Figure 5 shows typical examples of how the true covariances (available in the setting described above in this section) were estimated from the ensemble by (i) the localized ensemble (sample) covariances (used in the baseline stochastic EnKF method) and (ii) the model covariances generated by the fitted Locally Stationary Convolution Model. The left panel of Fig. 5 shows the variances and the right panel the correlations333 The “typical” realizations shown in Fig. 5 were selected as follows. Their single-realization mean absolute error (MAE) of the model variance (correlation) and its ratio to the respective MAE of the sample variance (localized sample correlation) were close to their counterparts computed by averaging over a large sample (of size 300). For correlations, MAE was computed by averaging over distances up to 15 mesh sizes..

### v5

@@ 5 LSEF: A Locally Stationary Ensemble Filter
The approach taken in this study is to use a spatial random field model to regularize the prior covariance matrix in sequential ensemble filtering. We call the resulting regularized EnKF the Locally Stationary Ensemble Filter (LSEF). Its forecast step coincides with that of the stochastic EnKF. The only difference between the LSEF’s and the traditional-EnKF’s analyses is how the proxy $\widehat{\bf B}$ to the true prior covariance matrix is defined. In this way, we can isolate the impact of our spatial-model-based regularization approach.
Testing and validation of our EnKF regularization technique can well be performed in the linear setting, avoiding unnecessary complications related to nonlinearities of observation and forecast models. For this reason, in the sequel we confine ourselves to the linear filtering setup. A reviews optimal and sub-optimal, ensemble and non-ensemble, time-discrete linear filters that are used in the numerical experiments presented below.
@@
In the LSEF analysis scheme, we postulate that the forecast error obeys the Locally Stationary Convolution Model. The model, or more specifically the spatial convolution kernel $u(x,\rho(x,y))$, is estimated from the forecast ensemble and a hyperprior distribution as detailed in section 4. With the estimated convolution kernel $\widehat{u}(x,\rho(x,y))$ in hand, we build the space-discrete approximation of Eq. (5):
@@
where $\boldsymbol{\xi}$ is the vector that represents the random field $\xi(x)$ evaluated on the spatial grid, $\boldsymbol{\alpha}\sim\mathcal{N}({\bf 0,I})$ and $\widehat{\bf W}$ is a matrix with the entries
$\widehat{w}_{ii^{\prime}}=\widehat{u}(x_{i},\rho(x_{i},x_{i^{\prime}}))\,\sqrt{\Delta x_{i^{\prime}}},$ (30)
where $\Delta x_{i^{\prime}}$ is the area of ${i^{\prime}}$th grid cell. Note that Eqs. (29) and (30) approximate the integral in Eq. (5) using the simplest rectangle rule, more accurate approximations will be considered elsewhere. Equation (29) entails that
@@
We note that the representation of the forecast-error covariance matrix $\widehat{\bf B}$ in the ‘square-root’ form, Eq. (31), is common in data assimilation practice because it provides efficient preconditioning of the analysis equations (e.g. Asch et al., 2016). The need for preconditioning follows from Eq. (41) because the matrix ${\bf B}^{-1}+{\bf H}^{\textsf{T}}{\bf R}^{-1}{\bf H}$ to be inverted is, normally, ill-conditioned. Substituting Eq. (31) into Eq. (41) and rearranging the terms yields
@@
Now the matrix to be inverted, ${\bf I}+\widehat{\bf W}^{\textsf{T}}{\bf H}^{\textsf{T}}{\bf R}^{-1}{\bf H}\widehat{\bf W}$, is well conditioned (as a sum of the identity matrix and a non-negative definite matrix). This is how the LSEF analysis is done, that is, without explicitly computing $\widehat{\bf B}$. The LSEF analysis assimilates all available observations at once, as this is done in variational data assimilation schemes (e.g. Asch et al., 2016).
@@ 6 Numerical experiments with static analyses
In this section, we experimentally examine the above LSEF analysis technique in the setting with known synthetic ‘truth’ (ground truth) that obeys the Locally Stationary Convolution Model on the sphere. Besides, the hyperprior distribution of the model parameters was known in these experiments as well, so the neural Bayes estimator could be trained using the perfect training sample. This setting allowed us to reveal the best possible performance our approach can provide.
We compared deterministic analyses produced by the following schemes:
  • True-B: the optimal analysis (which has access to the true ${\bf B}$ matrix).
  • LSEF-B: the new analysis proposed in this paper (with the neural Bayes estimator).
  • Mean-B: analysis with time-mean ${\bf B}$ (known in geosciences as 3D-Var).
  • EnKF-B: stochastic EnKF with tuned localization.
  • Hybrid-B: analysis with the half-sum of the time-mean ${\bf B}$ and EnkF’s ${\bf B}$ (known as EnVar).
See section 5 for the description of the LSEF analysis and A.3 and A.4 for the description of the other (reference) analyses. The quality of the analysis ensemble was not tested, it will be assessed in sequential filtering experiments presented in section 7.
@@ 6.1 Experimental methodology
Pseudo-random realizations of the ‘true’ spectral functions $\sigma_{\ell}(x)$ (which, we recall, are non-negative square roots of the local spectra $f_{\ell}(x)$) and the respective ‘true’ background error field $\xi(x)$ were generated following the Parametric Locally Stationary Convolution Model of ‘truth’ described in H. To give the reader an impression of the ground truth fields, we show in Fig. 1 two pseudo-random realizations444 We used SHTools (Wieczorek and Meschede, 2018) to generate pseudo-random fields on the sphere. of $\xi$ for two different values of the non-stationarity length parameter $\mu_{\rm NSL}$, which is defined in H.3 as the ratio of the non-stationarity length scale $\Lambda_{\rm NSL}$ and the median process length scale $L_{\rm med}$.
Figure 1: Two realizations of the random field produced by the parametric model of ‘truth’. Left: $\mu_{\rm NSL}=1$. Right: $\mu_{\rm NSL}=3$. x-axis: longitude index, y-axis: co-latitude index.
The left panel of Fig. 1 shows the field with the small $\mu_{\rm NSL}=1$, that is, with the strong spatial non-stationarity. The left panel of Fig. 1 displays a realization of the random field with the default (moderate) non-stationarity, $\mu_{\rm NSL}=3$. In both panels of Fig. 1, one can see a substantial degree of spatial non-stationarity, both in the field’s magnitude and length scale. With $\mu_{\rm NSL}=1$, Fig. 1(left), areas where the local length scale of $\xi$ is small (seen as patchy spots) and where the local length scale is large (smooth spots) are both seen to be smaller in size compared with those in the right panel. This reflects a smaller non-stationarity length scale with $\mu_{\rm NSL}=1$ than with $\mu_{\rm NSL}=3$, as expected.
@@
The true background-error vector $\boldsymbol{\xi}$ was computed by evaluating $\xi(\cdot){\mbox{$\,|\,$}}\sigma$ on the analysis grid (defined below in section 6.3). We assumed (without loss of generality) that ${\bf x}^{\rm f}=0$ so that the simulated ground truth was ${\bf x}={\bf x}^{\rm f}-\boldsymbol{\xi}=-\boldsymbol{\xi}$ and the ensemble members were ${\bf x}^{{\rm fe}(k)}={\bf x}^{\rm f}-\boldsymbol{\xi}^{(k)}=-\boldsymbol{\xi}^{(k)}$ (see A for notation).
@@
Having computed the true $\sigma_{\ell}(x)$, we used Eq. (6) to compute the true kernel $u(x,\rho)$. After that, we built the true ${\bf W}$ matrix following Eq. (30) and applied it to independent realizations of the white noise $\boldsymbol{\alpha}$ to generate the background-error vector $\boldsymbol{\xi}$ and the ensemble perturbations $\boldsymbol{\xi}^{(k)}$, see Eq. (29). The true covariance matrix (used by the optimal True-B analysis) was ${\bf B}={\bf W}{\bf W}^{\textsf{T}}$.
@@
With the truth ${\bf x}=-\boldsymbol{\xi}$ in hand, we generated observations ${\bf x}^{\rm obs}$ following Eq. (2), assimilated them using the five analysis techniques listed above, and compared the resulting deterministic analyses with the truth.
6.2 Details of the analysis schemes
The four sub-optimal analysis schemes differed only in the way the prior covariance matrix was specified. In the EnKF analysis, the localization matrix was specified using the popular Gaspari-Cohn correlation function (Gaspari and Cohn, 1999), whose length scale was tuned to obtain the best performance for each configuration of the model of truth and for each ensemble size separately.
In the Mean-B analysis, the static prior covariance matrix ${\bf B}_{\rm static}={\bf B}_{\rm mean}$ was computed for each configuration of the model of truth as follows. We generated $n_{\rm repl}=33$ replicates of the true spectral functions $\sigma_{\ell}(x)$. Given each replicate, $\sigma_{\ell}^{(i)}(x)$, we generated a sample of $n_{\rm sample}=10$ realizations of $\xi{\mbox{$\,|\,$}}\sigma$: $\xi^{(i,k)}(\cdot)$, where $i=1,2,\dots,n_{\rm repl}$ and $k=1,2,\dots,n_{\rm sample}$. Then, we expanded each realization in spherical harmonics, getting the spectral coefficients $\widetilde{\xi}^{(i,k)}_{\ell m}$. For each $\ell$, we summed $|\widetilde{\xi}^{(i,k)}_{\ell m}|^{2}$ over $m$ in the range from $-\ell$ to $\ell$, divided by $4\pi$, and averaged over $i$ and $k$, obtaining a sample estimate of the mean variance spectrum of $\xi$. From the mean variance spectrum, we computed the mean (stationary in space) covariance function and used it to build ${\bf B}_{\rm mean}$.
In the Hybrid-B analysis, a convex combination of the static and localized sample covariance matrices was used: $\widehat{\bf B}_{\rm hybr}=(1-w_{\rm e}){\bf B}_{\rm mean}+w_{\rm e}\widehat{\bf B}_{\rm EnKF}$, where $w_{\rm e}=0.5$ was found to be nearly optimal and used in all experiments.
@@ 6.2 Details of the analysis schemes
In LSEF-B, the bandpass filters’ spectral transfer functions were specified as
$H_{j}(\ell)=\exp\left(-\left|\frac{\ell-\ell^{\rm c}_{j}}{\Delta_{{\cal H}_{j}}}\right|^{q_{\rm shape}}\right),$ (33)
@@
where $\ell^{\rm c}_{j}$ is the central wavenumber of the $j$th filter, $\Delta_{{\cal H}_{j}}$ is the the half-width parameter, and $q_{\rm shape}$ is the shape parameter. In the experiments described below, we took ${q_{\rm shape}}=3$ (${q_{\rm shape}}=2$ also worked well). Both $\ell^{\rm c}_{j}$ and ${\Delta_{{\cal H}_{j}}}$ were taken exponentially growing with the band index $j$. Figure 2 shows the spectral transfer functions and the respective impulse response functions for six bandpass filters used in the experiments presented in this section. The spatial filtering was performed in spectral space.
@@ 6.3 Experimental setup
The default setup of the parametric model of truth is specified in H.4.
@@
Observations were located at randomly and independently sampled grid points selected with probabilities proportional to the spherical surfaces of the respective grid cells. The number of observations was $n_{\rm obs}=\frac{1}{2}n_{\rm x}$ (recall that $n_{\rm x}$ stands for the total number of grid points, that is, the length of the state vector ${\bf x}$). The observation-error variance was equal to the median background-error variance (we tried other choices and obtained qualitatively very similar results, not shown).
The ensemble size was between 5 and 80 with the default 20.
@@ 6.4 Training sample
Training the neural network involved in the estimation of the spatial model was done as follows. The Parametric Locally Stationary Convolution Model, see H, was used to generate $n_{\rm repl}=33$ replicates of the true spectral functions $\sigma_{\ell}(x)$. For each replicate of $\boldsymbol{\sigma}^{(i)}(\cdot)$ (with $i=1,2,\dots,n_{\rm repl}$), we generated an ensemble of $K$ pseudo-random realizations of $\xi{\mbox{$\,|\,$}}\sigma$, from which we computed vectors of sample band variances ${\bf d}^{(i)}(\cdot)$ at all analysis grid points following Eq. (21). As a result, with $n_{\rm x}$ grid points, we had the training sample of $n_{\rm train}=n_{\rm repl}\cdot n_{\rm x}$ pairs of vectors ${\bf d}^{(i)}(x)$ (inputs of the neural network) and $\boldsymbol{\sigma}^{(i)}(x)$ (targets to be predicted by the neural network). The neural network was trained only once, for the default setup.
@@ 6.5 Results
We compared the five deterministic analyses (listed at the beginning of section 6) in the default experimental setup and examined effects of varying the ensemble size $K$ and the hyperparameters $\varkappa$ (controls the strength of the non-stationarity, H.2) and $\mu_{\rm NSL}$ (controls the length scale of the non-stationarity, H.3).
Figure 3 shows dependencies of the RMSEs on the ensemble size. It is seen that all ensemble-based analyses improved with the increasing $K$, as it should be. The advantage of the LSEF-B analysis over the competitors was very significant.
@@
Figure 4 shows how the five analyses performed with the varying non-stationarity strength parameter $\varkappa$ (defined in H.2). The advantage of LSEF-B over EnKF-B was big and nearly uniform for all $\varkappa>1$ on the plot. It was Mean-B that suffered/benefited most from strong/weak non-stationarity, as expected. In the stationary regime (i.e., with $\varkappa=1$), Mean-B was almost as accurate as True-B because ${\bf B}_{\rm mean}$ was a very accurate estimate of the constant (in this regime) ${\bf B}_{\rm true}$. For this reason, Mean-B could not be rivaled by any ensemble-based analysis under stationarity.
@@
Finally in this section, we study the impact of the non-stationarity length scale parameter $\mu_{\rm NSL}$ (defined in H.3, see also section 6.1) on the performance of LSEF. Figure 5 displays RMSEs of the five analyses as functions of $\mu_{\rm NSL}$.
@@
Again, LSEF-B was significantly better than the other sub-optimal analyses, and that was the case even beyond the local stationarity assumption (section 3.6), which requires that the non-stationarity length scale is much larger than the typical length scale of the field itself. Indeed, LSEF-B was substantially better than the competing techniques even when $\mu_{\rm NSL}=1$ (see Fig. 5), that is, when the non-stationarity length scale $\Lambda_{\rm NSL}$ was equal to the median length scale $L_{\rm med}$ (see H.3).
It is important to notice that in LSEF-B, there was no need to retrain the neural network for different setups. This is in contrast to EnKF-B, where we had to tune the localization length scale for each triple $K,\varkappa,\mu_{\rm NSL}$. Similarly, ${\bf B}_{\rm mean}$ was computed for use in Mean-B and Hybrid-B for every pair $\varkappa,\mu_{\rm NSL}$. This suggests that the neural net-based LSEF-B was more robust than the competing analyses.
The neural network was trained on the sample of the same size as we used to calculate ${\bf B}_{\rm mean}$, yet it allowed LSEF-B to exhibit much better results than Hybrid-B. This suggests that the neural Bayes estimator is capable of selective blending of sample covariances with time-mean covariances. This is in contrast to Hybrid-B, which just mixes the two kinds of covariances with, effectively, equal weights for all spatial scales.
It is worth noting that the neural-Bayes-based LSEF-B analysis significantly outperformed an LSEF-B analysis based on the linear model estimator presented in section 4.4 (not shown). This was the expected outcome because the linear estimator (for which consistency is proven) is less accurate due to the linearity approximation.
6.6 Effects of covariance regularization on spatial covariances
We reproduced the above experimental methodology (the model of truth and the five analyses) on the circle and obtained very similar results (not shown). The code was written by the other author in a different programming language (R vs. Python).
Here we take advantage of the simpler circular domain to demonstrate how the regularization procedure proposed in this paper impacts spatial covariances. The spatial grid had 120 points and the ensemble size was 10.
Figure 6 shows typical examples of how the true covariances (available in the setting described above in this section) were estimated from the ensemble by (i) the localized sample covariances (used in the baseline stochastic EnKF method) and (ii) the model covariances generated by the estimated Locally Stationary Convolution Model. The left panel of Fig. 6 shows the variances and the right panel the correlations555 The realizations shown in Fig. 6 were typical in the following sense. Their single-realization mean absolute errors with respect to the truth were close (up to 10 percent difference) to the large-sample (of size 300) counterparts. For correlations, the mean absolute errors were computed by averaging over distances up to 15 mesh sizes..